1. You’re driving down the highway late one night at 20 m/s when a deer steps out onto the road 35 m in front on you. Your reaction time before stepping on the brakes is 0.50 seconds and the maximum deceleration of your car is 10 m/s2.

a) How much distance is between you and the deer when you come to a stop?

b) What is the maximum speed you could have and still not hit the deer?

2. Consider the two vectors A = 3 i – 2 j and B = - i -4 j.

Calculate the following: A + B, A – B, |A + B|, |A – B| and the directions (conventional angles) of A and B.
Answer #1

0 u= BOMIS 35 m t= 0.55 Let, so=distance covered = 2 deer t=0s speed = constant = 35 m/s Si=distance travelled = ut = 20x0.5=2. Here ☆ - 3ê -2 B²=-2 - 4 Atb A-B B | ATBL IA-BI between ASB o= angle ① Ā + B = 3ê ê -27_u - i - 6 ② Ā B = 3? - 2} + 2 + y}